If an electric device delivers a current of 5.0 A for 10 second show many electrons flow through the device?
Answer #1
Concepts and reason

Concepts used here will be charge, current and electron flow.

Initially from the definition of the current the net current flowing through the device is calculated and after that compares the current with the expression of the change in terms of the number of the electron.

Fundamentals

Current is defined as the rate of change of charge. The expression of the current is expresses as follows,

I=QtI = \frac{Q}{t}

Here QQis the associated charge and tt is the time related to flow.

The expression of the charger in terms of the number of the electron is expresses as follows,

Q=neQ = ne

Here nnis the number of electrons and eeis the charge per electron.

Rearranging gives,

I=netI = \frac{{ne}}{t}

And for n,

n=Iten = \frac{{It}}{e}

The expression of the current is as follows,

I=QtI = \frac{Q}{t}

The expression for the number of charge is expresses as follows,

Q=neQ = ne

Substitute neneforQQin the above expression of the current,

I=netI = \frac{{ne}}{t}

Rearrange the above expression in terms of the number of electron.

n=Iten = \frac{{It}}{e}

Calculate the number of electron.

Substitute 5.0A5.0{\rm{ A}}forII,10s10{\rm{ s}}fortt and 1.6×1019C1.6 \times {10^{ - 19}}{\rm{ C}}foreein the expression of the number of electron,

n=(5A)(10s)(1.6×1019C)=3.125×1020\begin{array}{c}\\n = \frac{{\left( {{\rm{5 A}}} \right)\left( {{\rm{10 s}}} \right)}}{{\left( {{\rm{1}}{\rm{.6 \times 1}}{{\rm{0}}^{ - {\rm{19}}}}{\rm{C}}} \right)}}\\\\ = 3.125 \times {10^{20}}\\\end{array}

Ans:

The required number of electrons flowing through the device is equal to3.125×10203.125 \times {10^{20}}.